OpenStudy (anonymous):

The function for the cost of materials to make a muffin is f(x) = Three fourths x + 3, where x is the number of muffins. The function for the selling price of those muffins is g(f(x)), where g(x) = 3x + 4. Find the selling price of 20 muffins.

ganeshie8 (ganeshie8):

\( \large f(x) = \frac{3}{4}x+3\) \(g(x) = 3x+4\) selling price = \(g(f(x)) = ?\)

ganeshie8 (ganeshie8):

u knw how to find/interpret \(g(f(x))\) ? :)

OpenStudy (anonymous):

No I don't.

ganeshie8 (ganeshie8):

its okay, u wil get it easily once u see it done

ganeshie8 (ganeshie8):

\(g(f(x))\) ^^ \(\ \large g(\frac{3}{4}x+3) \)

OpenStudy (anonymous):

Is it 18?

ganeshie8 (ganeshie8):

\(\large 3\left(\frac{3}{4}x+3\right) + 4\)

ganeshie8 (ganeshie8):

thats the g(f(x)) now plugin x=20 above

OpenStudy (anonymous):

Okay. Now I got 58..

ganeshie8 (ganeshie8):

58 = \(\large \color{red}{\checkmark}\)

OpenStudy (anonymous):

Thank you!

ganeshie8 (ganeshie8):

np :)